# Failure of uniform laws of large numbers for subdifferentials and beyond

### v1

@@ 2.3 Discussion
The previous subsections show that, for certain seemingly benign nonsmooth functions, obtaining a uniform LLN for subdifferentials is impossible. These negative results rule out a wide class of functions, but also shed light on where positive results might still exist. All our constructions involve functions that oscillate near some point; thus, at least in order to circumvent our counterexamples, some notion of tameness could be helpful [7].
Nevertheless, by considering weaker notions of convergence for subdifferentials, positive results can be obtained for a broad range of functions. One such notion is graphical convergence; see [12] for general results on graphical LLNs, [4] for convergence rates for weakly convex functions, and also [16] for results on monotone operators. Indeed, under suitable boundedness and joint measurability assumptions, it is shown in [12, Proposition 4.3] that the uniform law for enlarged subdifferentials in [19, Theorem 5] is equivalent to a graphical LLN on a compact set. Thus, our negative results can be interpreted as a separation between uniform and graphical LLNs for set-valued mappings.
We end with a concrete implication of this separation. For a random convex $f$, let $x_{0}$ be any $\varepsilon$-stationary point of $x\mapsto\tfrac{1}{\nu}\sum_{i=1}^{\nu}f(\bm{\xi}^{i},x)$; i.e., $\operatorname{dist}(0,\partial(\tfrac{1}{\nu}\sum_{i=1}^{\nu}f(\bm{\xi}^{i},\cdot))(x_{0}))\leq\varepsilon$. If the uniform LLN in (3) holds and $\nu$ is sufficiently large, then we can conclude that $x_{0}$ is a $2\varepsilon$-stationary point of $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$. In contrast, when only a graphical LLN holds, we can guarantee only the existence of some point $x_{0}^{\prime}$ near $x_{0}$ such that $x_{0}^{\prime}$ is $2\varepsilon$-stationary for $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$, and locating such an $x_{0}^{\prime}$ may be difficult even though its existence is assured.
@@ 3.1 A Probabilistic Gadget
which can be understood as the $k$th binary digit of $\xi-\lfloor\xi\rfloor$ using the terminating-zeros convention for dyadic rationals. For $\xi\in[0,1)$, we have the binary expansion
$\xi=\mathop{\sum}\nolimits_{k=1}^{\infty}2^{-k}\mathop{\textnormal{bit}}\nolimits_{k}(\xi).$
Let $\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be iid random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$, each with uniform distribution on $\Xi=[0,1]$. It is classical that the random variables $\{\mathop{\textnormal{bit}}\nolimits_{k}(\bm{\xi}^{i})\}_{i,k\in\mathbb{N}}$ are iid Bernoulli random variables with parameter $\tfrac{1}{2}$; see, e.g., [23, Section 4.6] and [8, Lemma 2.20].
@@
There exists an $\mathcal{F}$-measurable $E\subset\Omega$ with $\mathbb{P}(E)=1$ such that for any $\omega\in E$ and any $\nu\geq\bar{\bm{\nu}}(\omega)$, one has
$\mathop{\textnormal{bit}}\nolimits_{{\bm{k}^{\nu}(\omega)}}(\bm{\xi}^{1}(\omega))=\cdots=\mathop{\textnormal{bit}}\nolimits_{{\bm{k}^{\nu}(\omega)}}(\bm{\xi}^{\nu}(\omega))=1$
for some $\bar{\bm{\nu}}(\omega)\in\mathbb{N}$ and ${\bm{k}^{\nu}(\omega)}\leq\lceil 2^{\nu+1}\log(\nu+1)\rceil$.
@@
Let $K^{\nu}=\lceil 2^{\nu+1}\log(\nu+1)\rceil$. For any $\nu,k\in\mathbb{N}$, define sets
$E^{\nu}_{k}=\bigcap\nolimits_{i\in[\nu]}\{\mathop{\textnormal{bit}}\nolimits_{k}(\bm{\xi}^{i})=1\},\qquad E=\bigcup\nolimits_{\bar{\nu}\geq 1}\bigcap\nolimits_{\nu\geq\bar{\nu}}\bigcup\nolimits_{k\in[K^{\nu}]}E^{\nu}_{k},$
@@
$\mathbb{P}\Big(\bigcap\nolimits_{k\in[K^{\nu}]}(E^{\nu}_{k})^{\mathsf{c}}\Big)$ $=\prod\nolimits_{k\in[K^{\nu}]}\mathbb{P}((E^{\nu}_{k})^{\mathsf{c}})=\prod\nolimits_{k\in[K^{\nu}]}\Big(1-\prod\nolimits_{i=1}^{\nu}\mathbb{P}(\{\mathop{\textnormal{bit}}\nolimits_{k}(\bm{\xi}^{i})=1\})\Big)$
$=(1-2^{-\nu})^{K^{\nu}}\leq e^{-2^{-\nu}K^{\nu}}\leq\tfrac{1}{(\nu+1)^{2}}.$
Then, $\sum_{\nu=1}^{\infty}\mathbb{P}(\cap_{k\in[K^{\nu}]}(E^{\nu}_{k})^{\mathsf{c}})<\infty$. By Borel–Cantelli, we have
$\mathbb{P}\Big(\bigcap\nolimits_{\bar{\nu}\geq 1}\bigcup\nolimits_{\nu\geq\bar{\nu}}\bigcap\nolimits_{k\in[K^{\nu}]}(E^{\nu}_{k})^{\mathsf{c}}\Big)=\mathbb{P}(E^{\mathsf{c}})=0.$
For any $\nu\in\mathbb{N}$, define $\bm{k}^{\nu}(\omega)=1$ if $\omega\notin\cup_{k\in[K^{\nu}]}E_{k}^{\nu}$ and $\bar{\bm{\nu}}(\omega)=1$ if $\omega\notin E$. Otherwise, set
$\bm{k}^{\nu}(\omega)=\min\{k\in[K^{\nu}]\mid\omega\in E_{k}^{\nu}\},\quad\bar{\bm{\nu}}(\omega)=\min\{\bar{\nu}\geq 1\mid\omega\in\cap_{\nu\geq\bar{\nu}}\cup_{k\in[K^{\nu}]}E_{k}^{\nu}\},$
@@ 3.2 Proof of Theorem 1
Our proof is inspired by a construction that refutes one-sided uniform consistency in the setting of epigraphical LLNs; see [1, Example 4.1]. The key idea is to accumulate countably infinitely many affine pieces near the origin, with the slope of each piece determined by a random bit, $\mathop{\textnormal{bit}}\nolimits_{k}(\bm{\xi}^{i})\in\{0,1\}$. Then, with the help of Lemma 2, we can almost surely find a random point in $D^{\nu}$ at which the functions $\{f(\bm{\xi}^{i},\cdot)\mid i\in[\nu]\}$ are all $C^{1}$ with gradient consistently equal to $1$, thereby exhibiting a constant gap between the empirical and the average cases. Our proof is divided into five steps.
@@
which is well-defined as explained below. Define $\delta^{\nu}=(8\lceil 2^{\nu+1}\log(\nu+1)\rceil^{2})^{-1}.$
@@
The $\{0,1\}$-valued function $g(\xi,\cdot)$ is a sum of measurable indicator functions, then also measurable. Since $g(\xi,\cdot)$ is bounded and has only countably many points of discontinuity, the Lebesgue criterion implies that $g(\xi,\cdot)$ is Riemann integrable on compact sets, and hence $f$ is well defined. For any $x,y\in\mathbb{R}$, one has $|f(\xi,x)-f(\xi,y)|\leq|\int_{\min\{x,y\}}^{\max\{x,y\}}g(\xi,t)\textrm{d}t|\leq|x-y|$, hence $f(\xi,\cdot)$ is $1$-Lipschitz continuous. From Lemma 1, the function $\xi\mapsto\mathop{\textnormal{bit}}\nolimits_{k}(\xi)$ is measurable. By Tonelli, we have $f(\xi,x)=\sum_{k=1}^{\infty}\mathop{\textnormal{bit}}\nolimits_{k}(\xi)\int_{0}^{\max\{x,0\}}\mathds{1}_{B_{k}}\textrm{d}t<\infty$, hence $f(\cdot,x)$ is also measurable. Therefore, $f$ is Carathéodory. Moreover, for any $x\in\mathbb{R}$, $|\mathbb{E}[f(\bm{\xi},x)]|\leq\int_{0}^{\max\{x,0\}}1\textrm{d}t<\infty$.
@@
Hence, the function, $f(\xi,\cdot)$ is smooth on $\cup_{k}\mathbb{B}(p_{k},\Delta_{k})$. Moreover, for any $k\in\mathbb{N}$ and $y\in\mathbb{B}(p_{k},\Delta_{k})$, we can write $\nabla_{x}f(\xi,y)=g(\xi,y)=\mathop{\textnormal{bit}}\nolimits_{k}(\xi)$.
@@
Let $\mathcal{F}$-measurable $E\subset\Omega$ with $\mathbb{P}(E)=1$, $\bar{\bm{\nu}}$, and $\{\bm{k}^{\nu}\}_{\nu}$ be defined in Lemma 2. For any $\omega\in E$ and $\nu\geq\bar{\bm{\nu}}(\omega)$, one has $\mathop{\textnormal{bit}}\nolimits_{\bm{k}^{\nu}(\omega)}(\bm{\xi}^{i}(\omega))=1$ for all $i\in[\nu]$ and some $\bm{k}^{\nu}(\omega)\leq\lceil 2^{\nu+1}\log(\nu+1)\rceil$. Let $\bm{p}^{\nu}(\omega)=p_{\bm{k}^{\nu}(\omega)}\in\{p_{k}\}_{k}$. For any $y\in\mathbb{B}(\bm{p}^{\nu}(\omega),\Delta_{\bm{k}^{\nu}(\omega)})$, we have
@@
$0\leq\delta^{\nu}=\big(8\lceil 2^{\nu+1}\log(\nu+1)\rceil^{2}\big)^{-1}=\Delta_{\lceil 2^{\nu+1}\log(\nu+1)\rceil}\leq\Delta_{\bm{k}^{\nu}(\omega)}.$
Therefore, for any $\omega\in E$, we conclude that $\bm{p}^{\nu}(\omega)\in\{p_{k}\mid 1\leq k\leq\lceil 2^{\nu+1}\log(\nu+1)\rceil\}\subset D^{\nu}$ and

### v2

@@ 2.3 Discussion
The constructions used to prove Theorems 1 and 3 have implications beyond uniform LLNs. In this subsection, we discuss several of these consequences, focusing on SAA consistency and subdifferential approximation; see Table 1 for a summary of the results.
functions
$\mathbb{E}\partial$ and $\partial\mathbb{E}$
uniform
$\mathbb{E}\partial$
graphical
$\partial\mathbb{E}$
graphical
$\mathbb{E}\partial$
pointwise
$\partial\mathbb{E}$
pointwise
Lipschitz Corollary 1 Corollary 3 Corollary 4
Lip. & regular Corollary 2 [24, 16] [24, 16] [1] [1]
Table 1: Validity and failure of various laws of large numbers for subdifferentials. Uniform laws fail regardless of whether the expectation is taken before or after the subdifferential operator (see Remark 1), in contrast to graphical and pointwise laws; see Section 2.3.
Consistency of SAA.
The preceding subsections show that, for certain seemingly benign nonsmooth functions, a uniform LLN for subdifferentials cannot hold. This, however, does not preclude consistency of stationary points of the SAA objective, provided stationarity is interpreted in a weaker sense. Specifically, let a random vector $\bm{x}^{\nu}$ be a weak stationary point of the SAA objective $\tfrac{1}{\nu}\sum_{i=1}^{\nu}f(\bm{\xi}^{i},\cdot)$ $\mathbb{P}$-a.s.; i.e., $0\in\tfrac{1}{\nu}\sum_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},\bm{x}^{\nu})$. The qualifier “weak” reflects that, in general,
$\partial\big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\xi^{i},\cdot)\big)(x)\subset\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\xi^{i},x),\quad\partial\mathbb{E}[f(\bm{\xi},\cdot)](x)\subset\mathbb{E}[\partial_{x}f(\bm{\xi},x)],$
so every stationary point (i.e., $x$ such that $0\in\partial(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\xi^{i},\cdot))(x)$) is weak stationary, but not conversely. In the SAA setting of Section 1, if weak stationary points $\bm{x}^{\nu}\to x$ $\mathbb{P}$-a.s., then $0\in\mathbb{E}[\partial_{x}f(\bm{\xi},x)]$, i.e., $x$ is a weak stationary point of $\mathbb{E}[f(\bm{\xi},\cdot)]$; see, e.g., [30, Theorem 4.2].
Weak stationarity is used by many in stochastic programming (see, e.g., [30, 27, 6]), whereas stationarity is seldom used without assuming subdifferential regularity; see, e.g., [24, 13]. One may ask whether this is merely a matter of technique or reflects a genuine obstruction. The following corollary of our construction in Section 3.2 shows it is the latter.
Corollary 3 (consistency).
Under the assumptions of Theorem 1, and for $X$ and $\{\bm{\xi},\bm{\xi}^{\nu}\}_{\nu}$ constructed there, there exist a Carathéodory function $h:\Xi\times\mathbb{R}\to\mathbb{R}$ and random points $\{\bm{x}^{\nu}:\Omega\to X\}_{\nu}$ such that the following hold.
  1. (a)
    $\mathbb{P}$-a.s., for all sufficiently large $\nu\in\mathbb{N}$, $\bm{x}^{\nu}$ is a local minimizer of $\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}h(\bm{\xi}^{i},\cdot)$.
  2. (b)
    $\mathbb{P}$-a.s., one has $\bm{x}^{\nu}\to x\in X$, but $0\not\in\partial\mathbb{E}[h(\bm{\xi},\cdot)](x)$.
Corollary 3 shows that, as $\nu\to\infty$, cluster points of local minimizers of the SAA objective need not even be stationary. This strengthens the well-known observation that local optimality itself may fail to be preserved under SAA limits; see [3, (4.6)]. By the Fermat rule, a local minimizer is stationary. Therefore, Corollary 3 highlights the breakdown of a graphical LLN for subdifferentials when one considers $\partial\mathbb{E}[f(\bm{\xi},\cdot)]$ rather than $\mathbb{E}[\partial_{x}f(\bm{\xi},\cdot)]$; cf. [16, Example 4.4].
Subdifferential Approximation.
The subdifferential of an expectation function can be approximated in several ways, e.g., via smoothing [6]. Here we focus on the SAA approach (see, e.g., [4]), which is also tied to the Artstein–Vitale LLN for set-valued mappings [1]. Compared with the uniform and graphical laws discussed above, the Artstein–Vitale LLN can be viewed as a pointwise law. In the SAA setting of Section 1, it implies (see also [25, Theorem 7.53]) that,
$\lim_{\nu\to\infty}{d\kern-1.49994ptl}\Big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},x),\mathbb{E}[\partial_{x}f(\bm{\xi},x)]\Big)=0.$
As in the graphical laws of [24, 16], the above approximation is in the weak sense: it controls $\tfrac{1}{\nu}\sum_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},x)$ versus $\mathbb{E}[\partial_{x}f(\bm{\xi},x)]$, rather than $\partial(\tfrac{1}{\nu}\sum_{i=1}^{\nu}f(\bm{\xi}^{i},\cdot))(x)$ versus $\partial\mathbb{E}[f(\bm{\xi},\cdot)](x)$. It is thus natural to ask whether the latter type of convergence can ever hold in general. The following corollary of our construction in Section 3.2 gives a negative answer.
Corollary 4 (subdifferential approximation).
Under the assumptions of Theorem 1, and for $X$, $\{\bm{\xi},\bm{\xi}^{\nu}\}_{\nu}$, and $f$ constructed there, there exists $x\in X$ such that $\mathbb{P}$-a.s., one has
$\mathop{\rm liminf}_{\nu\to\infty}\operatorname{exs}\Big(\partial\big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\bm{\xi}^{i},\cdot)\big)(x);\partial\mathbb{E}[f(\bm{\xi},\cdot)](x)\Big)\geq\tfrac{1}{2}.$
Notably, the special pointwise case treated in Corollary 4 already suffices to rule out a variant of (3) when the expectation is taken inside the subdifferential operator. It does not, however, rule out the version with the expectation outside, since the corresponding pointwise statement always holds by the Artstein–Vitale LLN. By contrast, Theorem 1 shows that the uniform law (3) fails regardless of whether the expectation is placed inside or outside the subdifferential; see Remark 1.
DC Construction.
Analogues of Corollaries 3 and 4 also hold for a difference-of-convex (DC) variant of the construction used to prove Theorem 3; see Section 3.6 for details. Hence, restricting attention to DC functions with finitely many smooth pieces does not circumvent our negative results. In particular, the sufficient separation condition cannot be removed in [17, Corollary 5.1] while preserving consistency of d-stationary points.
@@ 3.1 A Probabilistic Gadget
which can be understood as the $k$th binary digit of $\xi$ using the terminating-zeros convention for dyadic rationals. For $\xi\in[0,1]$, we have the binary expansion $\xi=\mathop{\sum}\nolimits_{k=0}^{\infty}2^{-k}\mathop{\textnormal{bit}}\nolimits_{k}(\xi).$
Let $\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be iid random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$, each with uniform distribution on $\Xi=[0,1]$. It is classical that the random variables $\{\mathop{\textnormal{bit}}\nolimits_{k}(\bm{\xi}^{i})\}_{i,k\in\mathbb{N}}$ are iid Bernoulli random variables with parameter $\tfrac{1}{2}$; see, e.g., [29, Section 4.6] and [11, Lemma 2.20].
@@
There exist an $\mathcal{F}$-measurable set $E\subset\Omega$ with $\mathbb{P}(E)=1$ and $\mathbb{N}$-valued random variables $\bar{\bm{\nu}},\bm{k}^{1},\bm{k}^{2},\ldots$ such that for any $\omega\in E$ and any $\nu\geq\bar{\bm{\nu}}(\omega)$, one has
$\nu\leq{\bm{k}^{\nu}(\omega)}\leq\lceil 2^{\nu+1}\ln(\nu+1)\rceil\quad\text{and}\quad\mathop{\textnormal{bit}}\nolimits_{{\bm{k}^{\nu}(\omega)}}(\bm{\xi}^{1}(\omega))=\cdots=\mathop{\textnormal{bit}}\nolimits_{{\bm{k}^{\nu}(\omega)}}(\bm{\xi}^{\nu}(\omega))=1.$
@@
Let $K^{\nu}=\lceil 2^{\nu+1}\ln(\nu+1)\rceil$. For any $\nu,k\in\mathbb{N}$, define sets
$E^{\nu}_{k}=\bigcap\nolimits_{i\in[\nu]}\{\mathop{\textnormal{bit}}\nolimits_{k}(\bm{\xi}^{i})=1\},\qquad E=\bigcup\nolimits_{\bar{\nu}\geq 1}\bigcap\nolimits_{\nu\geq\bar{\nu}}\bigcup\nolimits_{k=\nu}^{K^{\nu}}E^{\nu}_{k},$
@@
$\mathbb{P}\Big(\bigcap\nolimits_{k=\nu}^{K^{\nu}}(E^{\nu}_{k})^{\mathsf{c}}\Big)$ $=\prod\nolimits_{k=\nu}^{K^{\nu}}\mathbb{P}((E^{\nu}_{k})^{\mathsf{c}})=\prod\nolimits_{k=\nu}^{K^{\nu}}\Big(1-\prod\nolimits_{i=1}^{\nu}\mathbb{P}(\{\mathop{\textnormal{bit}}\nolimits_{k}(\bm{\xi}^{i})=1\})\Big)$
$=(1-2^{-\nu})^{K^{\nu}-\nu+1}\leq e^{-2^{-\nu}(K^{\nu}-\nu+1)}\leq\tfrac{\sqrt{e}}{(\nu+1)^{2}}.$
Then, $\sum_{\nu=1}^{\infty}\mathbb{P}(\cap_{k=\nu}^{K^{\nu}}(E^{\nu}_{k})^{\mathsf{c}})<\infty$. By Borel–Cantelli, we have
$\mathbb{P}\Big(\bigcap\nolimits_{\bar{\nu}\geq 1}\bigcup\nolimits_{\nu\geq\bar{\nu}}\bigcap\nolimits_{k=\nu}^{K^{\nu}}(E^{\nu}_{k})^{\mathsf{c}}\Big)=\mathbb{P}(E^{\mathsf{c}})=0.$
For any $\nu\in\mathbb{N}$, define $\bm{k}^{\nu}(\omega)=\nu$ if $\omega\notin\cup_{k=\nu}^{K^{\nu}}E_{k}^{\nu}$ and $\bar{\bm{\nu}}(\omega)=1$ if $\omega\notin E$. Otherwise, set
$\bm{k}^{\nu}(\omega)=\min\{k\mid\nu\leq k\leq K^{\nu},\omega\in E_{k}^{\nu}\},\quad\bar{\bm{\nu}}(\omega)=\min\{\bar{\nu}\geq 1\mid\omega\in\cap_{\nu\geq\bar{\nu}}\cup_{k=\nu}^{K^{\nu}}E_{k}^{\nu}\},$
@@ 3.2 Proof of Theorem 1
Our proof is inspired by a construction that refutes one-sided uniform consistency of epigraphical LLNs; see [2, Example 4.1]. The key idea is to accumulate countably infinitely many affine pieces near the origin, with the slope of each piece determined by a random bit, $\mathop{\textnormal{bit}}\nolimits_{k}(\bm{\xi}^{i})\in\{0,1\}$. Then, using Lemma 2, we can almost surely find a random point in $D^{\nu}$ at which the functions $\{f(\bm{\xi}^{i},\cdot)\mid i\in[\nu]\}$ are all $C^{1}$ with gradient consistently equal to $1$, thereby exhibiting a constant gap between the empirical and the average cases. Our proof is divided into five steps.
@@
which is well-defined as explained below; see Figure 1. Define $\delta^{\nu}=(8\lceil 2^{\nu+1}\ln(\nu+1)\rceil^{2})^{-1}.$
@@
The $\{0,1\}$-valued function $g(\xi,\cdot)$ is a sum of measurable indicator functions, then also measurable. Since $g(\xi,\cdot)$ is bounded and has only countably many points of discontinuity, the Lebesgue criterion implies that $g(\xi,\cdot)$ is Riemann integrable on compact sets, and hence $f$ is well defined. For any $x,y\in\mathbb{R}$, one has $|f(\xi,x)-f(\xi,y)|\leq|\int_{\min\{x,y\}}^{\max\{x,y\}}g(\xi,t)\textrm{d}t|\leq|x-y|$, hence $f(\xi,\cdot)$ is $1$-Lipschitz continuous. From Lemma 1, the function $\xi\mapsto\mathop{\textnormal{bit}}\nolimits_{k}(\xi)$ is measurable. By Tonelli, we have $f(\xi,x)=\sum_{k=1}^{\infty}\mathop{\textnormal{bit}}\nolimits_{k}(\xi)\int_{0}^{\max\{x,0\}}\mathds{1}_{B_{k}}(t)\textrm{d}t<\infty$, hence $f(\cdot,x)$ is also measurable. Therefore, $f$ is Carathéodory. Moreover, for any $x\in\mathbb{R}$, $|\mathbb{E}[f(\bm{\xi},x)]|\leq\int_{0}^{\max\{x,0\}}1\textrm{d}t<\infty$.
@@
Hence, the function $f(\xi,\cdot)$ is smooth on $\cup_{k}\mathbb{B}(p_{k},\Delta_{k})$. Moreover, for any $k\in\mathbb{N}$ and $y\in\mathbb{B}(p_{k},\Delta_{k})$, we can write $\nabla_{x}f(\xi,y)=g(\xi,y)=\mathop{\textnormal{bit}}\nolimits_{k}(\xi)$.
@@
Let $\mathcal{F}$-measurable $E\subset\Omega$ with $\mathbb{P}(E)=1$, $\bar{\bm{\nu}}$, and $\{\bm{k}^{\nu}\}_{\nu}$ be defined in Lemma 2. For any $\omega\in E$ and $\nu\geq\bar{\bm{\nu}}(\omega)$, one has $\mathop{\textnormal{bit}}\nolimits_{\bm{k}^{\nu}(\omega)}(\bm{\xi}^{i}(\omega))=1$ for all $i\in[\nu]$ and some $\nu\leq\bm{k}^{\nu}(\omega)\leq\lceil 2^{\nu+1}\ln(\nu+1)\rceil$. Let $\bm{p}^{\nu}(\omega)=p_{\bm{k}^{\nu}(\omega)}\in\{p_{k}\}_{k}$. For any $y\in\mathbb{B}(\bm{p}^{\nu}(\omega),\Delta_{\bm{k}^{\nu}(\omega)})$, we have
@@
$0\leq\delta^{\nu}=\big(8\lceil 2^{\nu+1}\ln(\nu+1)\rceil^{2}\big)^{-1}=\Delta_{\lceil 2^{\nu+1}\ln(\nu+1)\rceil}\leq\Delta_{\bm{k}^{\nu}(\omega)}.$
Therefore, for any $\omega\in E$, we conclude that $\bm{p}^{\nu}(\omega)\in\{p_{k}\mid\nu\leq k\leq\lceil 2^{\nu+1}\ln(\nu+1)\rceil\}\subset D^{\nu}$ and